THE HIGH RESOLUTION INFRARED SPECTRUM OF THE LOWER WAVENUMBER BANDS OF ^{34}SO_{2}

\underline{W.J. LAFFERTY}, Optical Physics Laboratory, NIST, Gaithersburg, MD 20899, USA; J.-M.FLAUD, Laboratoire Inter Universitaire des Systemes Atmosphériques, CNRS, Université Paris 12, 61 Av du General de Gaulle, 94010 Créteil Cedex France; R.L. SAMS, Pacific Northwest Laboratory, P.O. Box 999, Richland, WA 99352, USA; EL HADJI ABIB NGOM, Univ. Cheikh Anta Diop, Ecole Supérieure Polytechnique. BP 5085, Dakar, Senegal.

The infrared spectrum of sulfur dioxide becomes of importance after volcanic eruptions in which large amounts of SO_{2} are emitted into the atmosphere. While the spectrum of the normal species, ^{32}SO_{2}, has been very well studied and an excellent set of molecular constants obtained, this is not the case for the rarer isotopic species, ^{34}SO_{2}, the abundance of which is about 4.2% of the normal species. In order to improve our knowledge of the ^{34}SO_{2} spectrum we have studied a sample of SO_{2} enriched to 95% in ^{34}S over a wavenumber region including all the fundamental bands, the relatively strong combination bands ν_{2}+ν_{3} and ν_{1}+ν_{3} as well as the weaker 2ν_{1}+ν_{3} and 2ν_{3} bands. The resolution was 0.0019 cm^{−1}. The two lowest hot bands of the bending mode, ν_{2}, have also been assigned as well as the hot bands of the other two fundamental bands and the ν_{1}+ν_{2}+ν_{3}-ν_{2} band. The spectroscopic constants obtained will be presented.